OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 My answer is 4 Am I right? :)

OpenStudy (uri):

How did you get to 4 @hunterhoff

OpenStudy (anonymous):

3x + 2y = 8 ---(1) 2x + 3y = 2 ---(2) Multiply equation (1) by 2 and equation (2) by 3 6x+4y=16 6x+9y=6 subtract -5y = 10 y=-2 plug this y into equation (1) 3x+2(-2) = 8 3x-4=8 3x=12 x=4

OpenStudy (uri):

3x + 2y = 8(multiply by 3) 2x + 3y = 2(multiply by-2) 9x+6y=24 -4x-6y=-4 ========= 5x=20 x=20/5 x=4 Do it this way..

OpenStudy (anonymous):

Ohh ok! Thats much easier, thanks :)